# TEAL: New Selection Strategy for Small Buffers in Experience Replay Class Incremenal Learning

### v1

@@ Abstract
Continual Learning is an unresolved challenge, whose relevance increases when considering modern applications. Unlike the human brain, trained deep neural networks suffer from a phenomenon called Catastrophic Forgetting, where they progressively lose previously acquired knowledge upon learning new tasks. To mitigate this problem, numerous methods have been developed, many relying on replaying past exemplars during new task training. However, as the memory allocated for replay decreases, the effectiveness of these approaches diminishes. On the other hand, maintaining a large memory for the purpose of replay is inefficient and often impractical. Here we introduce TEAL, a novel approach to populate the memory with exemplars, that can be integrated with various experience-replay methods and significantly enhance their performance on small memory buffers. We show that TEAL improves the average accuracy of the SOTA method XDER as well as ER and ER-ACE on several image recognition benchmarks, with a small memory buffer of 1-3 exemplars per class in the final task. This confirms the hypothesis that when memory is scarce, it is best to prioritize the most typical data.
@@ 1 Introduction
With the recent advances in deep neural networks, research interest in incremental learning has increased. Here, the problem of catastrophic forgetting, as defined by McCloskey and Cohen [20], can be severe. The necessity to incorporate new task knowledge into an already trained network has become more relevant given that training over extensive datasets is time-consuming. It is often impractical to retrain the network from scratch on the combined original and new task data. Furthermore, access to the already-trained data might be limited or non-existent.
Figure 1: Catastrophic forgetting when training CIFAR-10 [13] sequentially with only finetuning.
@@
Various methods have been developed to tackle this challenge across different frameworks, categorized by Van de Ven et al. [25] into three types: task-incremental, domain-incremental, and class-incremental learning. The core concept of incremental learning is that a model must learn tasks sequentially, one after another. Among these, class-incremental learning is considered the most challenging. In this scenario, each task introduces new classes, and the model must identify the specific class of each input without access to the task ID on which it was trained.
@@
There are several approaches to incremental learning, each with its own distinct assumptions and configurations. In this paper, we adhere to the constrained setup described by De Lange et al. [7], which does not rely on task boundaries during training or testing. This setup maintains a constant memory size throughout incremental training, ensuring that it does not exceed a predetermined limit established at the outset. At each incremental step, we can add new exemplars to the memory buffer only after space has been allocated for them.
@@
We focus our attention on a prevalent and rather successful framework called Experience Replay (ER), which involves storing a set of exemplars in memory and reusing them for rehearsal purposes while training on new tasks. Within this framework, several strategies exist for selecting which exemplars to retain in memory. Masana et al. [19] demonstrate that the most successful strategies are either random-sampling or Herding as defined by Welling [27]. Not surprisingly, the smaller the memory buffer is, the less effective the strategy is at mitigating catastrophic forgetting. This leads us to the following question: Are these strategies necessarily optimal for all memory sizes? In other words, could different strategies prove to be more effective depending on the size of the memory buffer?
@@
Our proposed method, TEAL, operates on similar principles, aiming to identify a set of representative exemplars that also exhibit diversity. An exemplar is deemed representative if its likelihood, when considered within the distribution of all points in an appropriate latent space, is high. Leveraging clustering ensures the selection of a diverse set of exemplars. This approach enhances the model’s capacity to retain previous knowledge while effectively learning new tasks, even when constrained by a small memory buffer. Central to the success of this approach is our ability to derive an appropriate latent space. The subsequent sections elaborate on our methodology, experimental setup, comparative performance against alternative selection strategies, and the enhancements TEAL offers to various existing class-incremental methods.
@@
Related work    Several approaches exist for incremental learning [see 7]. Alongside ER, alternatives include Generative Replay (GR) [23], Parameter Isolation, and Regularization. Similar to ER, GR replays data from previous tasks during new task training, but uses a generative model to create new samples instead of retaining exemplars seen by the model [6, 8, 9]. Parameter Isolation assigns distinct parameters to each task to reduce forgetting [18] by fixing parameters assigned to previous tasks, while Regularization-based methods [15] incorporate additional terms into the loss function to retain prior knowledge while learning from new data.
@@
These methods differ in how they utilize memory: ER stores exemplars, GR stores a generative model, and Parameter Isolation stores task-specific parameters. Regularization-based methods do not rely on memory. This diversity makes it challenging to compare methods directly. In particular, we note that since generative models tend to be very large, this approach is hardly suitable to the domain of small memory buffers addressed below. Hence, our focus in this paper is on ER methods only.
@@
Some CIL methods use an online learning setting, where each training example is seen only once. In this paper, we primarily use offline methods. Note that while few-shot incremental learning [24] may appear similar to our work on class-incremental learning (CIL) with a small memory buffer, there is a significant difference. We assume that the data for new tasks is initially sufficient for effective learning, whereas few-shot incremental learning inherently deals with a scarcity of labeled samples.
@@
Summary of contribution    We present TEAL, a novel strategy designed to select effective exemplars for small memory buffers. TEAL can be seamlessly integrated with replay-based class-incremental learning methods, significantly enhancing their performance.
@@ 2 Problem formulation
CIL involves learning from a sequence of disjoint tasks, where each task requires the classification of a set of classes, and every class belongs exclusively to a single task. At each incremental step, the learner receives the data of the new task and, in some cases, has access to a memory buffer containing exemplars from previous tasks. After each step, the learner is tested on the successful classification of all the classes it has seen, including those seen in previous steps.
More formally [see 19], a class-incremental learning problem $\mathcal{T}$ consists of a sequence of $T$ tasks. Each task $t$ contains labeled samples from $n_{t}$ different classes $X^{1},\dots,X^{n_{t}}$, where $X^{i}=\{(x_{1},i),\dots,(x_{m^{i}},i)\}$. The tasks do not share classes, i.e., $X^{i}\cap X^{j}=\emptyset~{}\forall i\neq j$. We denote by $N^{t}$ the total number of classes in all tasks up to task $t$: $N^{t}=\sum_{i=1}^{t}{n_{i}}$.
We consider the scenario where there is access to a memory buffer ${\mathcal{M}}$, which has a fixed size throughout the learning of $\mathcal{T}$. An incremental learner is a learning model (we consider only deep neural networks) that is trained sequentially on the tasks. Accordingly, the training on task $t$ is performed on data $X^{1},\dots,X^{n_{t}},{\mathcal{M}}$, where ${\mathcal{M}}$ contains exemplars from tasks $11,\dots,t-$ 1.
Different CIL methods may employ different loss functions, but at test time, they are all required to classify a given example into its predicted class while considering all previously seen classes.
@@ 3 Our method: TEAL
In active learning, it has been shown (empirically and formally) that when the number of labeled examples is small, it is best to choose the most typical examples for training. In the context of CIL and when the memory buffer is too small to truly represent the distribution of each class, we propose to adopt this strategy for the selection of points that are intended to populate the replay memory buffer. In other words, when the buffer is considerably small, it should contain representative exemplars and thus retain a more significant fraction of the previously acquired knowledge.
Our proposed method TEAL, Typicality Election Approach to continual Learning, aims to address only one component of the general Incremental Learning (IL) problem, namely, how to populate the memory buffer at the end of each IL iteration. Accordingly, after task $t$, TEAL should select for each seen class $n=\frac{|{\mathcal{M}}|}{N_{t}}$ exemplars to populate memory buffer ${\mathcal{M}}$.
Ideally, any new mechanism for memory population can be combined with any competitive Experience-Replay-Incremental-Learning (ER-IL) method in which this mechanism is a separate module. In accordance, we evaluate TEAL by considering alternative ER-IL methods, where we replace their native mechanism for buffer population by our proposed method. In almost all cases, this mechanism involves random sampling from the known labeled set, in order to faithfully represent the class distribution. In our evaluation study, we also consider an alternative population mechanism called Herding [22, 12].
3.1 Description of method
@@
To manage this, TEAL  maintains a list of the selected exemplars from each class ordered by typicality, reflecting the order in which they are to be deleted from ${\mathcal{M}}$. This allows the learner to remove the least typical points in the buffer, whenever new classes appear. To achieve this goal, whenever a new class is introduced, TEAL selects the initial set of representative exemplars through an iterative process. The selected list is ordered such that the most typical exemplars, which are to be retained for as long as possible, are positioned first, while those that will be removed first are positioned last.
More specifically, in each repeat of this iterative selection process and for each class separately, TEAL selects a small number of exemplars (a fraction of $n$, where $n$ denotes the number of exemplars assigned to the respective class). Let $S_{i}$ denote the set of exemplars after the $i^{th}$ iteration, for a total of $k$ iterations. Note that $S_{1}\subseteq S_{2}\subseteq\dots\subseteq S_{k}$ and that $|S_{k}|=n$. The sizes $|S_{1}|,\dots,|S_{k}|$ define the selection pace, indicating the pace of selecting exemplars from new class data.
@@
To initialize the selection process, we need a suitable embedding space for the class exemplars. To this end we train a deep model on all the available training data, and use activations in its penultimate layer as a representation for the new classes [cf. 22]. Subsequently, the selection process in the $i^{th}$ iteration involves 2 steps (see pseudocode in Alg. 1):
Step 1: Clustering
@@
Since typical examples tend to cluster in bunches [10], and in order to ensure diversity, we seek typical exemplars from different regions of the embedding space. We achieve this by dividing the set of labeled points into $|S_{i}|$ clusters, where $|S_{i}|$ represents the number of exemplars that should be selected from this class after the $i^{th}$ iteration.
Step 2: Typicality
@@
We then select the most typical point from each of the $|S_{i}|-|S_{i-1}|$ largest uncovered clusters, where an uncovered cluster is a cluster from which no point has already been selected. We order the selected points by their typicality score.
  Input: $X^{c}$eqparboxCOMMENT# a set of exemplars from class $c$
  Input: $n$eqparboxCOMMENT# number of examples to choose
@@
  Require: $\phi:{\mathcal{X}}\to{\mathbb{R}}^{d}$eqparboxCOMMENT# current feature function
  Require: $s_{1},\dots,s_{k}$eqparboxCOMMENT# iterations pace, $s_{1}\leq\dots\leq s_{k}=n$
  Output: ${\mathcal{M}}_{c}$eqparboxCOMMENT# a set of n exemplars chosen from $X^{c}$
  $X^{c}_{emb}\leftarrow\phi(X^{c})$
  ${\mathcal{M}}_{c}\leftarrow\emptyset$
  for all $i=1,\dots,k$ do
     $C_{1},\dots,C_{s_{i}}\leftarrow$ clustering_algorithm$(X^{c}_{emb},\ s_{i})$eqparboxCOMMENT# $|C_{1}|\geq\dots\geq|C_{s_{i}}|$
     for all $j=1,\dots,s_{i}-s_{i-1}$ do
        add $\underset{x\in C_{j}}{\mathrm{argmax}}\,\{Typicality(x)\}$ to ${\mathcal{M}}_{c}$
     end for
  end for
  return ${\mathcal{M}}_{c}$
Algorithm 1 TEALConstructExemplarSet
3.2 Using TEAL for incremental learning
@@
The integration of TEAL into a general class-incremental algorithm is described in Algs. 2-3.
  Input: ${\mathcal{M}}_{c}=(m_{1},\dots,m_{|{\mathcal{M}}_{c}|})$eqparboxCOMMENT# current exemplars set
  Input: $n$eqparboxCOMMENT# number of examples to keep
  return $(m_{1},\dots,m_{n})$
Algorithm 2 ReduceExemplarSet
  Input: $X^{t}=\{X^{N^{t-1}+1},\dots,X^{N^{t}}\}$eqparboxCOMMENT# training examples in per-class sets
  Require: $\theta^{t-1}$eqparboxCOMMENT# current model parameters
@@
  Require: ${\mathcal{M}}=({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t-1}})$eqparboxCOMMENT# current exemplar sets
  $\theta^{t}\leftarrow\textrm{training\_model}(X^{t},{\mathcal{M}};\theta^{t-1})$
  $n\leftarrow\frac{|{\mathcal{M}}|}{N^{t}}$eqparboxCOMMENT# number of exemplars per class
  for $c\in\{1,\dots,N^{t-1}\}$ do
     ${\mathcal{M}}_{c}\leftarrow ReduceExemplarSet({\mathcal{M}}_{c},n)$
  end for
  for $c\in\{N^{t-1}+1,\dots,N^{t}\}$ do
     ${\mathcal{M}}_{c}\leftarrow ConstructExemplarSet(X^{c},n,\theta^{t})$
  end for
  ${\mathcal{M}}\leftarrow({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t}})$
Algorithm 3 IncrementalTrain

### v2

# TEAL: New Selection Strategy for Small Buffers in Experience Replay Class Incremental Learning

@@ Abstract
Continual Learning is an unresolved challenge, whose relevance increases when considering modern applications. Unlike the human brain, trained deep neural networks suffer from a phenomenon called catastrophic forgetting, wherein they progressively lose previously acquired knowledge upon learning new tasks. To mitigate this problem, numerous methods have been developed, many relying on the replay of past exemplars during new task training. However, as the memory allocated for replay decreases, the effectiveness of these approaches diminishes. On the other hand, maintaining a large memory for the purpose of replay is inefficient and often impractical. Here we introduce TEAL, a novel approach to populate the memory with exemplars, that can be integrated with various experience-replay methods and significantly enhance their performance with small memory buffers. We show that TEAL enhances the average accuracy of existing class-incremental methods and outperforms other selection strategies, achieving state-of-the-art performance even with small memory buffers of 1-3 exemplars per class in the final task. This confirms our initial hypothesis that when memory is scarce, it is best to prioritize the most typical data. Code is available at this https URL.
@@ 1 Introduction
With the recent advances in deep neural networks, there has been a growing research interest in incremental learning. The need to integrate new task knowledge into an already trained network has become increasingly important, especially considering the time-consuming nature of training on large datasets. Retraining the network from scratch on both the original and new task data is often impractical, and access to the original training data may be limited or unavailable. In this context, catastrophic forgetting, as described by McCloskey and Cohen (1989), can be particularly severe.
@@
To address this challenge, various methods have been developed across different frameworks. Van de Ven, Tuytelaars, and Tolias (2022) categorizes these methods into three types: task-incremental, domain-incremental, and class-incremental learning. The fundamental idea behind incremental learning is that a model must sequentially learn tasks, one after the other. Among these approaches, Class-Incremental Learning (CIL) is recognized as the most challenging. Here, each task introduces new classes, and the model must accurately identify the class of each input without access to the corresponding task ID, see Fig. 1.
Figure 1: Illustration of CIL with Experience Replay.
@@
Incremental learning can be approached in various ways, each with its own assumptions and configurations. In this paper, we follow the constrained setup outlined by De Lange et al. (2021), which does not depend on task boundaries during either training or testing. This approach maintains a fixed memory size throughout incremental training, ensuring it stays within a predefined limit set from the beginning. At each incremental step, new exemplars can be added to the memory buffer only after sufficient space has been vacated.
@@
We focus our attention on a prevalent and rather successful framework called Experience Replay (ER), which involves storing a set of exemplars in memory and reusing them for rehearsal purposes while training on new tasks. Within this framework, several strategies exist for selecting which exemplars to retain in memory. Not surprisingly, the smaller the memory buffer is, the less effective the strategy is at mitigating catastrophic forgetting. This leads us to the following question: Are these strategies necessarily optimal for all memory sizes? In other words, can different strategies be found suitable for different sizes of the memory buffer?
In active learning, it has been shown (empirically and formally) that when the number of labeled examples is small, it is best to choose the most typical examples for training (Hacohen, Dekel, and Weinshall 2022). In the context of CIL and when the memory buffer is too small to truly represent the distribution of each class, we propose to adopt this strategy for the selection of points that are intended to populate the replay memory buffer. In other words, when the buffer is considerably small, it should contain representative exemplars and thus retain a more significant fraction of the previously acquired knowledge.
Our proposed method TEAL, Typicality Election Approach to continual Learning, is primarily targeted at scenarios with small buffers. While the use of small buffers may seem too strict, there are situations where maintaining a large memory for replay is simply not feasible. For example, applications running on mobile devices face severe memory constraints and could greatly benefit from methods specifically designed for small buffers.
@@
Accordingly, TEAL  aims to identify a set of representative exemplars that also exhibit diversity. An exemplar is deemed representative if its likelihood, when considered within the distribution of all points, is high. To ensure diversity, data clustering is leveraged. Central to the success of our approach is the ability to derive an appropriate latent space, where data clustering and likelihood estimation can be reliably obtained. Ideally, any mechanism for memory population can be combined with any competitive Experience-Replay-Incremental-Learning (ER-IL) method in which this mechanism is a separate module. In accordance, we evaluate our approach in Section 4 by considering alternative ER-IL methods, where we replace their native mechanism for buffer population by TEAL. The method, methodology and its experimental evaluation are described in the rest of this paper, with emphasis on the enhancements TEAL offers to various existing class-incremental methods.
@@
Related work    Several approaches exist for incremental learning (see De Lange et al. 2021), including Experience Replay (ER), Generative Replay (GR) (Shin et al. 2017), Parameter Isolation, and Regularization. Similar to ER, GR replays data from previous tasks during new task training, but uses a generative model to create new samples instead of retaining exemplars seen by the model (Choi, El-Khamy, and Lee 2021; Gao and Liu 2023; Gautam et al. 2024). Parameter Isolation assigns distinct parameters to each task to reduce forgetting (Mallya and Lazebnik 2018) by fixing parameters assigned to previous tasks, while Regularization-based methods (Li and Hoiem 2017) incorporate additional terms into the loss function to retain prior knowledge while learning from new data.
@@
These methods differ in the ways they utilize memory: ER stores exemplars, GR stores a generative model, and Parameter Isolation stores task-specific parameters. Regularization-based methods do not rely on memory. This diversity makes it challenging to compare methods directly. In particular, we note that since generative models tend to be very large, GR is hardly suitable to the domain of small memory buffers addressed here. A similar concern can be raised concerning Parameter Isolation methods. Hence, our focus in this paper is on ER.
@@
Note that while few-shot incremental learning (Tian et al. 2024) may appear similar to our work on CIL with a small memory buffer, there is a significant difference. Specifically, we assume that the data for new tasks is initially sufficient for effective learning, whereas few-shot incremental learning inherently deals with a scarcity of labeled samples.
When comparing strategies for the population of the memory buffer, Masana et al. (2022) demonstrate that the most successful strategies are either random-sampling or Herding as defined by Welling (2009). The latter strategy involves retaining a set of exemplars whose mean is closest to the class mean. Another successful strategy presented by Bang et al. (2021) leverages classification uncertainty and data augmentation to enhance the diversity of data instances (Uncertainty). Other selection strategies, such as GSS (Aljundi et al. 2019) and selecting the exemplars with the highest entropy of the softmax outputs (Chaudhry et al. 2018a), have been shown in previous studies to be inferior to random sampling and Herding (Masana et al. 2022; Prabhu et al. 2023). Therefore, we do not include them in our comparative empirical evaluation.
Recently, Hacohen and Tuytelaars (2024) proposed a selection strategy called Goldilocks, which retains exemplars learned at an intermediate pace, accommodating various buffer sizes. Since the results presented in this paper are primarily within a task-incremental framework (which yields higher average accuracy) and the paper does not provide code, we do not include this method in our comparisons.
@@
Summary of contribution    We present TEAL, a novel strategy designed to select effective exemplars for small memory buffers. TEAL is seamlessly integrated with a number of SOTA replay-based class-incremental learning method, and significantly enhances their performance.
@@ 2 Problem formulation
A CIL problem $\mathcal{T}$ consists of a sequence of $T$ tasks. Each task $t\in T$ contains a set of classes $C^{t}=(c^{1},\dots,c^{n_{t}})$ and labeled samples from these classes $X_{t}=(X^{1},\dots,X^{n_{t}})$, where $X^{i}=\{(x_{1},i),\dots,(x_{m^{i}},i)\}$. The tasks do not share classes, i.e., $C^{t_{i}}\cap C^{t_{j}}=\emptyset~{}\forall i\neq j$. We denote by $N^{t}$ the total number of classes in all tasks up to task $t$: $N^{t}=\sum_{i=1}^{t}{n_{i}}$.
We consider the scenario where there is access to a memory buffer ${\mathcal{M}}$, which has a fixed size throughout the learning of $\mathcal{T}$. An incremental learner is a learning model (we consider only deep neural networks) that is trained sequentially on the tasks. Accordingly, the training on task $t$ is performed on data $X_{t}\cup{\mathcal{M}}$, where ${\mathcal{M}}$ contains stored exemplars from tasks $11,\dots,t-$ 1. At test time, a CIL method is required to classify a given example into its predicted class while considering all previously seen classes.
Our proposed method TEAL is described in Section 3. It aims to address only one component of the general Incremental Learning (IL) problem, namely, how to populate the memory buffer at the end of each IL iteration. Accordingly, after task $t$, TEAL should select for each seen class $n=\frac{|{\mathcal{M}}|}{N_{t}}$ exemplars to populate memory buffer ${\mathcal{M}}$.
@@ 3 Our method: TEAL
Figure 2: Illustration of TEAL’s iterative class selection process, which establishes a priority order for the selected set. Initially, an embedding space is generated separately for each class (shown on the left). Samples are then selected iteratively with $s_{1}=4$, $s_{2}=7$, and $s_{3}=9$ (see text for details). Each row on the right panel represents one iteration: the left image displays the $s_{i}$ clusters (obtained using K-means) with the previously selected points $S_{i-1}$ marked with ’o’, while the right image shows the updated set $S_{i}$, with newly selected samples marked with ’X’. In the last iteration 3 clusters remain uncovered, but only 2 samples are selected, leaving the red cluster uncovered by $S_{3}$.
@@
To address this challenge, TEAL maintains a priority list of selected exemplars from each class, which reflects the order in which they should be removed from ${\mathcal{M}}$. This approach allows the learner to remove the least typical points from the buffer when new classes emerge. The list is structured so that the most typical and diverse exemplars, which should be retained as long as possible, are positioned at the top, while those slated for earlier removal are positioned at the bottom.
More specifically, consider a fixed class, and let $n$ denote the number of exemplars assigned to that class. TEAL repeatedly selects a small fraction of $n$, generating a sequence of subsets $S_{1}\subseteq S_{2}\subseteq\dots\subseteq S_{k}$, where $s_{i}=|S_{i}|$ and $s_{1}\leq\dots\leq s_{k}=n$. The sizes ${s_{i}}$ define the selection pace, indicating the rate at which exemplars are accumulated. The inclusion relation induces the priority order: points in $S_{k}\setminus S_{k-1}$ are removed first, while points in $S_{1}$ are removed last. This process is illustrated in Fig. 2.
@@
To initialize the selection process, we need a suitable embedding space for the class exemplars. To this end we train a deep model on all the available training data, and use activations in its penultimate layer as a representation for the new classes. Subsequently, the selection process in the $i^{th}$ iteration involves 2 steps (see pseudocode in Alg. 1):
Step 1: Clustering.
@@
In order to ensure diversity, we seek typical exemplars from different regions of the embedding space. When constructing set $S_{i}$, we achieve this by dividing the set of labeled points into $s_{i}$ clusters using K-Means (Lloyd 1982).
Step 2: Typicality.
@@
We then select the most typical point from each of the $s_{i}-s_{i-1}$ largest uncovered clusters, where an uncovered cluster is a cluster from which no point has already been selected.
Algorithm 1 ConstructExemplarSet
Input: a set of exemplars from class $c$ $X^{c}$, number of examples to choose $n$
@@
Require: current feature function $\phi:{\mathcal{X}}\to{\mathbb{R}}^{d}$, iterations pace $s_{1},\dots,s_{k}=n$
Output: a set of n exemplars ${\mathcal{M}}_{c}\subseteq X^{c}$
1:  $X^{c}_{emb}\leftarrow\phi(X^{c})$
2:  ${\mathcal{M}}_{c}\leftarrow\emptyset$
3:  for all $i=1,\dots,k$ do
4:     $C_{1},\dots,C_{s_{i}}\leftarrow$ clustering_algorithm$(X^{c}_{emb},\ s_{i})$triangleright $|C_{1}|\geq\dots\geq|C_{s_{i}}|$
5:     for all $j=1,\dots,s_{i}-s_{i-1}$ do
6:        if $C_{j}$ is uncovered  then
7:           add $\operatorname*{argmax}_{x\in C_{j}}\{Typicality(x)\}$ to ${\mathcal{M}}_{c}$
8:        end if
9:     end for
10:  end for
11:  return ${\mathcal{M}}_{c}$
@@
The integration of TEAL into a general class-incremental algorithm is described in Alg. 2.
Algorithm 2 IncrementalTraining
Input: training examples $X_{t}$
@@
Require: current exemplar sets ${\mathcal{M}}=({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t-1}})$, current model parameters $\theta^{t-1}$
1:  $\theta^{t}\leftarrow\textrm{training\_model}(X_{t},{\mathcal{M}};\theta^{t-1})$
2:  $n\leftarrow\frac{|{\mathcal{M}}|}{N^{t}}$
3:  for $c\in\{1,\dots,N^{t-1}\}$ do
4:     ${\mathcal{M}}_{c}\leftarrow$ first $n$ exemplars in ${\mathcal{M}}_{c}$
5:  end for
6:  for $c\in\{N^{t-1}+1,\dots,N^{t}\}$ do
7:     ${\mathcal{M}}_{c}\leftarrow ConstructExemplarSet(X^{c},n,\theta^{t})$
8:  end for
9:  ${\mathcal{M}}\leftarrow({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t}})$